A charge is located at the origin, another charge is located somewhere on the +y axis, another charge is located somewhere on the -y axis, and a fourth is located somewhere on the +x axis. The angle of the total electric force on the charge at the origin is 27 degrees counter-clockwise from the +x axis. Then, the charge at the origin changes sign and its magnitude (absolute value of its charge) increases by a factor of 3.7. All other charges do not move or change their signs or magnitudes. What is the angle of the total electric force on the charge at the origin now measured in degrees counter-clockwise from the +x axis?
Answer #1

for the net force on the first charge to be in the first quadrant , the y-component of the net force must be along positive Y-axis and X-component of the net force must be along the positive x-axis. hence the forces look like this

from the vector diagram

tan27 = (F2 - F3)/F4

F2 F2- F3 F2 F3 F4 F4 27 F4 Fs F4 F4 F4 F2-F3 F2

when the charge at origin is made 3.7 times and its sign change , the magnitude of each force becomes 3.7 times and the direction of each force reverse.

so we get the diagram as

tanheta = (F'2 - F'3)/F'4

where

F'2 = 3.7 F2

F'3 = 3.7 F3

F'4 = 3.7 F4

tanheta = ((3.7)F2 - (3.7)F3)/((3.7)F4 )

tanheta =(F2 - F3)/F4

tanheta = tan27

heta = 27

phi = angle from positive x-axis counterclockwise = 180 + heta = 180 + 27 = 207 deg